16. You will flip a fair coin 200 times. What is the approximate probability you will get at least 53% Heads?
(A) 0.8485 (B) 0.1711 (C) 0.8023 (D) 0.1977 (E) 0.1515
Answer #1

Solution:

Given:  a fair coin is flipped 200 times. Thus n = 200 and p = probability of head = 0.5

We have to find the  approximate probability you will get at least 53% Heads.

That is we have to find:

Plp > 0.53) = ....

Find z score:

р-р Грх(1-р)

2 = 0.53 -0.5 0.5x(1-0.5) 200

0.03 0.5x0.5 200

0.03 0.00125

0.03 0.035355

=0.85

Thus we get:
P p > 0.53) = P(Z > 0.85)

P p > 0.53) = 1- P(Z < 0.85)

Look in z table for z = 0.8 and 0.05 and find area.

「 | 0.0 lol 10.2 10.3 10.4 10.5 10.6 |o.7 10.8 .00 .01 .02 .03 .04 1.05 - 106 5000 5040 .5080 .5120 5160 499 5239 5398 .5438

P( Z < 0.85) =0.8023

Thus

P p > 0.53) = 1- P(Z < 0.85)

P p > 0.53) = 1 - 0.8023

P p > 0.53) = 0.1977

Thus  the approximate probability you will get at least 53% Heads is 0.1977
Answer #1

Solution:

Given:  a fair coin is flipped 200 times. Thus n = 200 and p = probability of head = 0.5

We have to find the  approximate probability you will get at least 53% Heads.

That is we have to find:

Plp > 0.53) = ....

Find z score:

р-р Грх(1-р)

2 = 0.53 -0.5 0.5x(1-0.5) 200

0.03 0.5x0.5 200

0.03 0.00125

0.03 0.035355

=0.85

Thus we get:
P p > 0.53) = P(Z > 0.85)

P p > 0.53) = 1- P(Z < 0.85)

Look in z table for z = 0.8 and 0.05 and find area.

「 | 0.0 lol 10.2 10.3 10.4 10.5 10.6 |o.7 10.8 .00 .01 .02 .03 .04 1.05 - 106 5000 5040 .5080 .5120 5160 499 5239 5398 .5438

P( Z < 0.85) =0.8023

Thus

P p > 0.53) = 1- P(Z < 0.85)

P p > 0.53) = 1 - 0.8023

P p > 0.53) = 0.1977

Thus  the approximate probability you will get at least 53% Heads is 0.1977
Answer #1

Solution:

Given:  a fair coin is flipped 200 times. Thus n = 200 and p = probability of head = 0.5

We have to find the  approximate probability you will get at least 53% Heads.

That is we have to find:

Plp > 0.53) = ....

Find z score:

р-р Грх(1-р)

2 = 0.53 -0.5 0.5x(1-0.5) 200

0.03 0.5x0.5 200

0.03 0.00125

0.03 0.035355

=0.85

Thus we get:
P p > 0.53) = P(Z > 0.85)

P p > 0.53) = 1- P(Z < 0.85)

Look in z table for z = 0.8 and 0.05 and find area.

「 | 0.0 lol 10.2 10.3 10.4 10.5 10.6 |o.7 10.8 .00 .01 .02 .03 .04 1.05 - 106 5000 5040 .5080 .5120 5160 499 5239 5398 .5438

P( Z < 0.85) =0.8023

Thus

P p > 0.53) = 1- P(Z < 0.85)

P p > 0.53) = 1 - 0.8023

P p > 0.53) = 0.1977

Thus  the approximate probability you will get at least 53% Heads is 0.1977